OpenStudy (anonymous):

Find the area of the triangle with a base of one and four fifths and height of five sevenths. I came up with an answer but want to confirm.

OpenStudy (whpalmer4):

What is your answer?

OpenStudy (anonymous):

My answer is 9/14 yrd squared

OpenStudy (whpalmer4):

\[A = \frac{1}{2}bh= \frac{1}{2}*1\frac{4}{5}*\frac{5}{7} = \frac{1}{2}*\frac{9}{5}*\frac{5}{7} = \frac{9}{14}\checkmark\]